OpenStudy (anonymous):

Which expression has a value of –12? A. –4 – (–8) B. 3 – 9 C. –16 – (–4) D. –6 – 18

OpenStudy (anonymous):

D?

OpenStudy (♪chibiterasu):

No. A negative minus a positive is just increasing the negative.

OpenStudy (anonymous):

C

OpenStudy (♪chibiterasu):

Yes.

OpenStudy (bishoppatton):

It is definitely C.

OpenStudy (♪chibiterasu):

-16 -(-4) is the same as -16 + 4, which is -12.

OpenStudy (bishoppatton):

yes... she is right

OpenStudy (anonymous):

so negative and negative is a positive, positive and positive is positive

OpenStudy (bishoppatton):

yes

OpenStudy (anonymous):

what about positive and negative and negative and positive

OpenStudy (anonymous):

i get mixed up on those

OpenStudy (♪chibiterasu):

16 - 4 = 12 -16 + 4 = -12

OpenStudy (♪chibiterasu):

-16 + 4 = decreaseing the negative

OpenStudy (♪chibiterasu):